OpenStudy (anonymous):

using the provided rule, determine if the sequence is arithmetic or geometric and generate the first 4 terms. an= 2n-3

OpenStudy (anonymous):

a1= 2.1 -3 = -1 a2=2.2-3 = +1 a3=2.3-3= 3 a4=2.4-3=5 here a2-a1=2 a3-a2=2 a4-a3=2 hence Arithmetic sequence...